OpenStudy (anonymous):

Consider the line whose equation is 4x – 8y = 8. Write the equation of the line, in point-slope form, that is perpendicular to the line and passes through the point (1, 2).

OpenStudy (anonymous):

Point slope form is defined as y - y1 = m(x - x1)

OpenStudy (anonymous):

Find the slope which is y2-y1/x2-x1

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

Plug your slope back into the equation

OpenStudy (anonymous):

y=y1=1/2(x-x1)

OpenStudy (anonymous):

I find it easier to use standard form which is y = mx+b

OpenStudy (anonymous):

m = slope

OpenStudy (anonymous):

@satellite73 help i am confused

OpenStudy (anonymous):

We will start over... have you found y?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Rewrite the equation in standard form as y =

OpenStudy (anonymous):

2=1/2x-1?

OpenStudy (anonymous):

No... it should be in the form y = mx+b

OpenStudy (anonymous):

y=1/2x-1

OpenStudy (anonymous):

Here is how u want to approach this. I can't tell you are close, just little mistakes here and there I sounds like.

OpenStudy (anonymous):

That is what I was trying to explain. Thanks Demo

OpenStudy (anonymous):

this is the question right Write the equation of the line, in point-slope form, that is perpendicular to the line and passes through the point (1, 2).?

OpenStudy (anonymous):

yss

OpenStudy (anonymous):

did you get the answer, or do you still need help?

OpenStudy (anonymous):

still need help

OpenStudy (anonymous):

ok but you did most of it already you found the slope of \[4x – 8y = 8.\] is \(m=\frac{1}{2}\)

OpenStudy (anonymous):

the perpendicular line has a slope that is the "negative reciprocal" i.e. flip it and change the sign the negative reciprocal of \(\frac{1}{2}\) is \(-2\)

OpenStudy (anonymous):

now the question really becomes find the equation of the line with slope \(-2\) through \((1, 2)\)

OpenStudy (anonymous):

for that you use the "point slope" formula \[y-y_1=m(x-x_1)\] which in our case is \[y-2=-2(x-1)\]

OpenStudy (anonymous):

to put it in the "slope intercept" form, solve for \(y\) the steps are always the same first multiply out on the right and get \[y-2=-2x+2\] then add \(2\) to finish with \[y=-2x+4\] and that is done

OpenStudy (anonymous):

ok thanks i know get it

OpenStudy (anonymous):

good not too hard you got another one to try?

OpenStudy (anonymous):

i have a differnt one on how to graph a eqation

OpenStudy (anonymous):

like the last one?

OpenStudy (anonymous):

um like the last one we graphed kinda x+3y<6 that is a greater than or equal sign

OpenStudy (anonymous):

\[x+3y\leq 6\] i.e. "less than or equal to" right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok lets do this the easy way

OpenStudy (anonymous):

if \(x=0\) you get \(3y=6\) or \(y=2\) so we know \((0,2)\) is on the graph, where the line hits the \(y\) axis

OpenStudy (anonymous):

if \(y=0\) you get \(x=6\) so you know \((6,0)\) is on the graph, where the line crosses the \(x\) axis lets plot those two points only

OpenStudy (anonymous):



OpenStudy (anonymous):

now since it is \(x+3y\leq 6\) the graph is not just a line but the whole region BELOW the line

OpenStudy (anonymous):

shade in all stuff underneath

OpenStudy (anonymous):

ok so thats it

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

yw